# MythTriage: Scalable Detection of Opioid Use Disorder Myths on a Video-Sharing Platform

### v1

@@ Abstract
Understanding the prevalence of misinformation in health topics online can inform public health policies and interventions. However, measuring such misinformation at scale remains a challenge, particularly for high-stakes but understudied topics like opioid-use disorder (OUD)—a leading cause of death in the U.S. We present the first large-scale study of OUD-related myths on YouTube, a widely-used platform for health information. With clinical experts, we validate 8 pervasive myths and release an expert-labeled video dataset. To scale labeling, we introduce MythTriage, an efficient triage pipeline that uses a lightweight model for routine cases and defers harder ones to a high-performing, but costlier, large language model (LLM). MythTriage achieves up to 0.86 macro F1-score while estimated to reduce annotation time and financial cost by over 76% compared to experts and full LLM labeling. We analyze 2.9K search results and 343K recommendations, uncovering how myths persist on YouTube and offering actionable insights for public health and platform moderation. [^1] We will release the datasets and models upon acceptance.
Warning: Some content of this paper, included to contextualize our data, are misleading.
@@
Hayoung Jung1        Shravika Mittal2        Ananya Aatreya1 Navreet Kaur1    Munmun De Choudhury2    Tanushree Mitra1 1University of Washington  2Georgia Institute of Technology {hjung10, tmitra}@uw.edu
Figure 1: Study Overview. (1.) We curated opioid-related topics and queries (1.1), then collected YouTube search and recommendation results (1.2). (2.) To label myths, clinical experts validated 8 myths (with examples shown in the orange box), refined the annotation guidelines, and provided 2.5K labels (2.1). We compare three potential labeling approaches: experts labeling, LLM labeling, and MythTriage—a scalable pipeline using lightweight distilled models for easy cases and defers hard ones to high-performing, but costly LLM. (3.) Using MythTriage, we analyzed the labeled dataset at scale, offering actionable insights for platform moderation and public health.
@@ 1 Introduction
Online platforms are a key source of health information Dubey et al. (2014), with video-sharing platforms like YouTube playing an increasingly prominent role in shaping public understanding during public health crises Bora et al. (2018); Khatri et al. (2020). However, online platforms are also a conduit for widespread misinformation that can undermine public health efforts Milmo (2022). A particular instance is the case of opioid use disorder (OUD) – a leading cause of death in the U.S, with 108K drug overdose deaths in 2022 NIDA (2023). Facing offline stigma, individuals with OUD often rely on online platforms for health information and recovery guidance Balsamo et al. (2023). But online myths—e.g., medication for addiction treatment (MAT) is simply replacing one drug with another—fuel treatment hesitancy, distrust in healthcare, and stigma ElSherief et al. (2021a); Woo et al. (2017). This has undermined efforts to promote clinically-approved MAT NASEM (2019).
Understanding the scale and spread of such misinformation is crucial for health officials and platforms seeking to design effective interventions Juneja and Mitra (2021). While prior works have acknowledged this gap and explored social dynamics and discourse in online health communities Bunting et al. (2021); Chancellor et al. (2019), large-scale analyses of the OUD-related myth prevalence, especially on video-sharing platforms, remain limited. Detecting misinformation on video platforms at scale remains challenging, as it requires domain expertise and intensive labeling of large volumes of content. While recent works highlight the potential of large language models (LLMs) to address this scale challenge in social science research Ranjit et al. (2024); Dammu et al. (2024), their increasing compute demands and high API inference cost—especially on long-form video content—limits their widespread use for large-scale misinformation detection, particularly in high-stakes health issues.
To address these gaps, we present the first large-scale study of OUD-related myths on YouTube, illustrated in Figure 1. We construct two datasets: OUD Search Dataset of 2.9K search results (1.8K unique videos) from 73 trending queries across four opioid and four treatment topics, and OUD Recommendation Dataset of 343K recommendations (164K unique videos) obtained by crawling the top-four recommendations per unique video in OUD Search Dataset, going five levels deep. In collaboration with clinical experts, we validate 8 pervasive myths (see Table 8 for list of myths and examples), refine the annotation guidelines, and construct a gold-standard dataset of 310 videos labeled across all myths, totaling 2.5K expert labels.
@@ 2 Related Works
Stigma and Misinformation in High-Stakes Health Contexts. Prior works have investigated online platforms and LLMs for stigma and misinformation in high-stakes health contexts Jung et al. (2025); Nguyen et al. (2024); Kaur et al. (2024), with efforts to employ LLMs to reduce stigma and support well-being Song et al. (2025); Mittal et al. (2025). Among health conditions, OUD is among the most stigmatized, often viewed as a result of “willful choice” rather than a chronic, treatable disease Corrigan (2017); Kennedy-Hendricks et al. (2017). Such a narrative fuels persistent myths that undermine harm reduction and MAT Garett and Young (2022); Kruis et al. (2021). To escape offline stigma, many turn to online recovery communities for support and information D’Agostino et al. (2017); Balsamo et al. (2023), but these spaces also contain harmful OUD myths. A few studies have quantified OUD-related myths online Mittal et al. (2024); ElSherief et al. (2021b, 2024b), but these efforts have been limited to a small set of myths and text-based platforms like Reddit and Twitter. Building on these efforts, we collaborate with clinical experts and introduce MythTriage, a scalable pipeline for detecting 8 distinct OUD-related myths across large video corpora. This work presents the first large-scale empirical analysis of OUD-related myth prevalence on YouTube, a challenging task that requires collaboration with clinical experts.
@@ 3.1.1 Selecting Topics
To identify opioid topics, we used a curated lexicon of 156 keywords covering opioid drugs, medication-assisted treatments (MAT), and prescription medicines. This lexicon—developed in consultation with public health experts and prior literature in ElSherief et al. (2021a)—includes generic names (e.g., Oxycodone), trade names (e.g., OxyContin), and street names (e.g., OC) to ensure comprehensive coverage of opioid-related topics.
Since collecting data for all 156 keywords is impractical, we used Google Trends (Trends) to identify the four most popular opioid and four most popular MAT keywords, yielding eight keywords in total. Trends reflects real-world search interests and suggests related queries. We systematically filtered out overly broad keywords (e.g., “Water”), those lacking Trends data, or those with fewer than five related queries, reducing the set to 28 keywords.
To rank keywords by popularity, we conducted pairwise comparisons of keywords in Trends. [^2] Trends limits comparisons to five keywords and lacks an API, so we used SerpApi (2024) to automate the comparisons. For each pair, Trends returns relative interest scores (ranging from 0 to 100); the higher-scoring keyword is considered more popular and thus “wins.” To rank keywords, we computed win-rates across all pairs (see Table 5). To capture the most popular topics across opioids and MAT, we selected the top 4 opioid and MAT keywords, [^3] While not MAT, we include “Kratom” due to debunked claims of its efficacy in treating OUD Mayo Clinic (2024). with the highest win-rates, yielding 8 final topics (Table 1). §A details Trends configuration and pairwise comparison.
@@ 3.2 Collecting Data on YouTube
With the topics and queries, we collected YouTube search and recommendation results to measure the prevalence of OUD-related myths.
@@ 3.2.1 Collecting Search Results
For each video, we gathered metadata, including title, description, transcript, tags, and its rank in the results, creating the OUD Search Dataset with 2,893 search results (1,776 unique videos).
@@ 3.2.2 Collecting Recommendation Results
To measure the prevalence of myths in YouTube recommendations, we used a cascaded approach following Albadi et al. (2022) to gather recommendations from December 18th to 20th, 2024. Using Google’s InnerTube API Bulled (2025), we collected the top four recommended videos per unique video in the OUD Search Dataset, as in Albadi et al. (2022). This yielded 6,356 level 1 video recommendations (3,107 unique videos). We repeated this process through level 5, collecting the top four recommendations per unique video at the previous level, resulting in a total of 342,707 recommendation links and 164,085 unique videos across all levels (see Table 7). We retrieved the metadata for these videos through the YouTube Data API to create the OUD Recommendation Dataset. [^4] We could not retrieve data for 274 videos (0.2% of data).
@@ 4.1 How do we know what is a myth?
To identify myths, we drew from prior literature and clinical sources. ElSherief et al. (2024a) conducted a systematic review of four online platforms, where three public health experts identified five prevalent OUD-related myths based on substance use literature Heimer et al. (2019); Wakeman and Barnett (2018). We supplemented these with three additional myths from clinical sources, debunking pervasive myths about MAT National Addiction Specialist (2023); SCTN (2024); Hill (2023); Peckham (2022). All selected myths are recognized by major health organizations, such as the U.S. SAMHSA (2023) and Johns Hopkins Medicine (2020), and were validated by clinical researchers, as described below. Table 2 lists the 8 myths.
@@ 4.2 Expert-Annotated Gold Standard Dataset
Creating the Gold Standard Dataset. With the sampled videos, we conducted multiple rounds of annotations with six clinical researchers as experts using our annotation scheme (§B.1), resulting in a gold-standard dataset of 2,480 high-quality labels (8 myths×$\times$310 videos). We detail annotation process in §B.3. Experts reported an average annotation time of 3 minutes per video.
@@ 5 Labeling for Myths in YouTube Videos
With 164K videos in the OUD Recommendation Dataset, manual annotation by experts is infeasible. To scale the labeling of myths, we leverage LLMs for myth detection (§5.1), distill lightweight classifiers (§5.2), and implement MythTriage, an efficient triage pipeline in which the lightweight classifiers route challenging cases to the LLM (§5.3). We apply MythTriage to the OUD Recommendation Dataset, evaluating its efficiency and cost effectiveness (§5.4).
@@ 5.2 Distillation for Myth Detection
While GPT-4o performs well, its financial and computational costs make the model impractical for large-scale labeling of our task. Meanwhile, we lack sufficient high-quality, expert-labeled data to fine-tune a model (Table 11). To address these challenges, we use GPT-4o, the best-performing LLM for our task, to generate high-quality synthetic labels Zheng et al. (2023) for the 1,466 videos in OUD Search Dataset that were not annotated by experts, and train a lightweight student model Rao et al. (2023); Park et al. (2024) on this synthetic data. This approach minimizes API and computational costs and avoids the instability of relying on proprietary LLMs, whose behaviors can change over time OpenAI (2024b), while achieving strong performance on our gold-standard dataset. For each myth, we train DeBERTa-v3-base He et al. (2021) with training and experimental details in §D.
Results. Table 3 shows that DeBERTa-v3-base, trained on GPT-4o-generated synthetic labels, achieves macro F1-scores between 0.60 and 0.78 across myths, with scores $\geq 0.75$ on four myths. These results demonstrate strong performance despite the models’ smaller size and usage of synthetic data. The results highlight the models’ effectiveness on a high-stake, complex video classification task and their suitability for large-scale labeling. Full evaluation metrics are in Table 14, with additional details in §D.3.
@@ 5.3 MythTriage Implementation
With GPT-4o offering stronger performance and the distilled model enabling efficient large-scale labeling, MythTriage can combine their strengths—using the lightweight model for routine cases and deferring harder ones to the strong, but costly LLM to optimize both cost and performance. To decide which examples to defer, we use two strategies: (1) Maximum Softmax Probability (MSP), which uses the predicted class softmax probability as a simple, effective proxy for model confidence Lakshminarayanan et al. (2017), deferring examples below a chosen MSP threshold, and (2) Validation Error Tendencies (VET), which defers predictions from classes with low validation performance (e.g., class-specific F1 < 0.8). We also evaluated combining MSP and VET. We discuss alternative deferral strategies, threshold selection method, and results motivating MSP and VET in §E.
@@ 7.1 Prevalence of Myths in Search Results
Kratom has the highest prevalence of myths across topics (Figure 2), with 36% of search results supporting and only 22% opposing OUD-related myths. This is concerning given the widespread but debunked claims about Kratom’s effectiveness in treating OUD. The high prevalence of myth-supporting content may mislead users towards unsafe alternatives Mayo Clinic (2024), undermining evidence-based treatments like MAT. While Heroin and Methadone have slightly negative bias scores (-0.02 and -0.04), they still show high levels of myth-supporting content (22% and 36%). These findings can help platforms prioritize moderation on high-risk topics and inform public health officials where misinformation is most concentrated.
@@
Switching from the default “Relevance” filter increases exposure to myths. In Figure 3, the “Relevance” filter consistently returns search results that lean towards opposing myths (bias scores from -0.04 to -0.26). In contrast, alternative filters—“Upload Date,” View Count,” and “Rating”—consistently shift bias scores in a more positive direction, indicating increased prevalence of myth. This finding aligns with a prior work in the domain of COVID-19 misinformation Jung et al. (2025), and suggests that users seeking recent, most-viewed, or highly-rated videos (more likes than dislikes) are more likely to encounter myths. This is concerning, as users may place more trust in popular content and those seeking the latest information about OUD may be at greater risk of encountering myths. These findings present an opportunity for YouTube to improve moderation by enhancing safeguards in non-default search filters.

### v2

@@ Abstract
Understanding the prevalence of misinformation in health topics online can inform public health policies and interventions. However, measuring such misinformation at scale remains a challenge, particularly for high-stakes but understudied topics like opioid-use disorder (OUD)—a leading cause of death in the U.S. We present the first large-scale study of OUD-related myths on YouTube, a widely-used platform for health information. With clinical experts, we validate 8 pervasive myths and release an expert-labeled video dataset. To scale labeling, we introduce MythTriage, an efficient triage pipeline that uses a lightweight model for routine cases and defers harder ones to a high-performing, but costlier, large language model (LLM). MythTriage achieves up to 0.86 macro F1-score while estimated to reduce annotation time and financial cost by over 76% compared to experts and full LLM labeling. We analyze 2.9K search results and 343K recommendations, uncovering how myths persist on YouTube and offering actionable insights for public health and platform moderation. [^1] Code and Data: https://github.com/hayoungjungg/MythTriage, Models: https://huggingface.co/SocialCompUW/youtube-opioid-myth-detect-M1
Warning: Some content of this paper, included to contextualize our data, is misleading.
@@
Hayoung Jung1thanks: Most work done at the University of Washington.        Shravika Mittal3        Ananya Aatreya2 Navreet Kaur2    Munmun De Choudhury3    Tanushree Mitra2 1Princeton University 2University of Washington 3Georgia Institute of Technology Correspondence: hayoung@cs.princeton.edu, tmitra@uw.edu
Figure 1: Study Overview. (1.) We curated opioid-related topics and queries (1.1), then collected YouTube search and recommendation results (1.2). (2.) To label myths, clinical experts validated 8 myths (with examples shown in the orange box), refined the annotation guidelines, and provided 2.5K labels (2.1). We compare three potential labeling approaches: expert labeling, LLM labeling, and MythTriage—a scalable pipeline using lightweight distilled models for easy cases and defers hard ones to high-performing, but costly LLM. (3.) Using MythTriage, we analyzed the labeled dataset at scale, offering actionable insights for platform moderation and public health.
@@ 1 Introduction
Online platforms are a key source of health information Dubey et al. (2014), with video-sharing platforms like YouTube playing an increasingly prominent role in shaping public understanding during public health crises Bora et al. (2018); Khatri et al. (2020). However, online platforms are also a conduit for widespread misinformation that can undermine public health efforts Milmo (2022). A particular instance is the case of opioid use disorder (OUD) – a leading cause of death in the U.S, with 108K drug overdose deaths in 2022 NIDA (2023). Facing offline stigma, individuals with OUD often rely on online platforms for health information and recovery guidance Balsamo et al. (2023). But online myths—e.g., medication for addiction treatment (MAT) is simply replacing one drug with another—fuel treatment hesitancy, distrust in healthcare, and stigma ElSherief et al. (2021); Woo et al. (2017). This has undermined efforts to promote clinically-approved MAT NASEM (2019).
Understanding the scale and spread of such misinformation is crucial for health officials and platforms seeking to design effective interventions Juneja and Mitra (2021). While prior works have acknowledged this gap and explored social dynamics and discourse in online health communities Bunting et al. (2021); Chancellor et al. (2019), large-scale analyses of the OUD-related myth prevalence, especially on video-sharing platforms, remain limited. Detecting misinformation on video platforms at scale remains challenging, as it requires domain expertise and intensive labeling of large volumes of content. While recent works highlight the potential of large language models (LLMs) to address this scale challenge in social science research Ranjit et al. (2024); Dammu et al. (2024), their increasing compute demands and high API inference cost—especially on long-form video content—limit their widespread use for large-scale misinformation detection, particularly in high-stakes health issues.
To address these gaps, we present the first large-scale study of OUD-related myths on YouTube, illustrated in Figure 1. We construct two datasets: OUD Search Dataset of 2.9K search results (1.8K unique videos) from 73 trending queries across four opioid and four treatment topics, and OUD Recommendation Dataset of 343K recommendations (164K unique videos) obtained by crawling the top-four recommendations per unique video in OUD Search Dataset, going five levels deep. In collaboration with clinical experts, we validate 8 pervasive myths (see Table 9 for the list of myths and examples), refine the annotation guidelines, and construct a gold-standard dataset of 310 videos labeled across all myths, totaling 2.5K expert labels.
@@ 2 Related Works
In contrast to prior model cascading frameworks evaluated mainly on standard benchmarks such as CoLA Varshney and Baral (2022), SQuAD Mamou et al. (2022), or CIFAR-10 Berestizshevsky and Even (2019), our work demonstrates the practical value of these frameworks by integrating and validating them in a real-world, high-stakes health contexts, contributing a scalable, extensible pipeline for large-scale OUD myth detection in video-sharing platforms.
Stigma and Misinformation in High-Stakes Health Contexts. Prior works have investigated online platforms and LLMs for stigma and misinformation in high-stakes health contexts Jung et al. (2025); Nguyen et al. (2024); Kaur et al. (2024), with efforts to employ LLMs to reduce stigma and support well-being Song et al. (2025); Mittal et al. (2025). Among health conditions, OUD is among the most stigmatized, often viewed as a result of “willful choice” rather than a chronic, treatable disease Corrigan (2017); Kennedy-Hendricks et al. (2017). Such a narrative fuels persistent myths that undermine harm reduction and MAT Garett and Young (2022); Kruis et al. (2021). To escape offline stigma, many turn to online recovery communities for support and information D’Agostino et al. (2017); Balsamo et al. (2023), but these spaces also contain harmful OUD myths. A few studies have quantified OUD-related myths online Mittal et al. (2024); ElSherief et al. (2021, 2024), but these efforts have been limited to a small set of myths and text-based platforms like Reddit and Twitter. Building on these efforts, we collaborate with clinical experts and introduce MythTriage, a scalable pipeline for detecting 8 distinct OUD-related myths across large video corpora. This work presents the first large-scale empirical analysis of OUD-related myth prevalence on YouTube, a challenging task that requires collaboration with clinical experts.
@@ 3.1.1 Selecting Topics
To identify opioid topics, we used a curated lexicon of 156 keywords covering opioid drugs, medication-assisted treatments (MAT), and prescription medicines. This lexicon—developed in consultation with public health experts and prior literature in ElSherief et al. (2021)—includes generic names (e.g., Oxycodone), trade names (e.g., OxyContin), and street names (e.g., OC) to ensure comprehensive coverage of opioid-related topics.
Since collecting data for all 156 keywords is impractical, we used Google Trends (Trends) to identify the four most popular opioid and four most popular MAT keywords, yielding eight keywords in total. Trends reflects real-world search interests and suggests related queries. We systematically filtered out overly broad keywords (e.g., “Water”), those lacking Trends data, or those with fewer than five related queries, filtering the set to 28 keywords (see Table 6). The complete set of the 28 keywords is shown in Appendix Table 6.
Using the set of 28 keywords, we conducted pairwise comparisons of keywords in Trends to rank them by popularity. [^2] Trends limits comparisons to five keywords and lacks an API, so we used SerpApi (2024) to automate the comparisons. For each pair, Trends returns relative interest scores (ranging from 0 to 100); the higher-scoring keyword is considered more popular and thus “wins.” We computed win-rates across all pairs (see Table 6). To identify the most popular topics across opioids and MAT, we selected the top 4 opioid and MAT keywords, [^3] While not MAT, we include “Kratom” due to debunked claims of its efficacy in treating OUD Mayo Clinic (2024). with the highest win-rates from the set of 28 keywords, yielding 8 final topics (Table 1). §A details Trends configuration and pairwise comparison.
@@ 3.1.2 Selecting Search Queries
Dataset Total Unique Expert Synthetic
OUD Search 2,893 1,776 310 1,466
OUD Recommendation 342,707 164,085 0 164,085
Table 2: Summary of the OUD Search and OUD Recommendation Dataset, showing the number of total videos (Total), unique videos (Unique), expert-annotated videos (Expert), and GPT-4o-labeled videos (Synthetic). Expert annotations were focused on the OUD Search Dataset to build a gold-standard dataset (§4.2). Both datasets require the same video labeling, but the OUD Recommendation Dataset is much larger (164K videos) and noisier, making expert annotation infeasible; we therefore used MythTriage for efficient labeling.
@@ 3.2 Collecting Data on YouTube
With the topics and queries, we collected YouTube search and recommendation results to measure the prevalence of OUD-related myths. Table 2 summarizes the datasets, reporting counts of total, unique, expert-annotated, and GPT-4o-labeled videos.
@@ 3.2.1 Collecting Search Results
For each video, we gathered metadata, including title, description, transcript, tags, and its rank in the results, creating the OUD Search Dataset with 2,893 search results (1,776 unique videos; see Table 2).
@@ 3.2.2 Collecting Recommendation Results
To measure the prevalence of myths in YouTube recommendations, we used a cascaded approach following Albadi et al. (2022) to gather recommendations from December 18th to 20th, 2024. Using Google’s InnerTube API Bulled (2025), we collected the top four recommended videos per unique video in the OUD Search Dataset, as in Albadi et al. (2022). This yielded 6,356 level 1 video recommendations (3,107 unique videos). We repeated this process through level 5, collecting the top four recommendations per unique video at the previous level, resulting in a total of 342,707 recommendation links and 164,085 unique videos across all levels (Table 2). Detailed counts of videos per recommendation levels are provided in Table 8. We retrieved the metadata for these videos through the YouTube Data API to create the OUD Recommendation Dataset. [^4] We could not retrieve data for 274 videos (0.2% of data).
@@ 4.1 How do we know what is a myth?
To identify myths, we drew from prior literature and clinical sources. ElSherief et al. (2024) conducted a systematic review of four online platforms, where three public health experts identified five prevalent OUD-related myths based on substance use literature Heimer et al. (2019); Wakeman and Barnett (2018). We supplemented these with three additional myths from clinical sources, debunking pervasive myths about MAT National Addiction Specialist (2023); SCTN (2024); Hill (2023); Peckham (2022). All selected myths are recognized by major health organizations, such as the U.S. SAMHSA (2023) and Johns Hopkins Medicine (2020), and were validated by clinical researchers, as described below. Table 3 lists the 8 myths.
@@ 4.2 Expert-Annotated Gold Standard Dataset
[…]
@@ 5 Labeling for Myths in YouTube Videos
With 164K videos in the OUD Recommendation Dataset, manual annotation by experts is infeasible. To scale the labeling of the OUD-related myths, we leverage LLMs for myth detection (§5.1), distill lightweight classifiers (§5.2), and implement MythTriage, an efficient triage pipeline in which the lightweight classifiers route challenging cases to the LLM (§5.3). We apply MythTriage to the OUD Recommendation Dataset, evaluating its efficiency and cost effectiveness (§5.4).
@@ 5.2 Distillation for Myth Detection
While GPT-4o performs well, its financial and computational costs make the model impractical for large-scale labeling of our task. Meanwhile, we lack sufficient high-quality, expert-labeled data to fine-tune a model (Table 12). To address these challenges, we adopt a distillation approach, as prior works have shown that student models can be effectively trained from high-performing teacher models Rao et al. (2023); Park et al. (2024); Zheng et al. (2023). As discussed in §5.1, GPT-4o achieved the strongest performance on our task (macro F1: 0.82-0.87 across myths), making it a suitable teacher model. We therefore use GPT-4o to generate high-quality synthetic labels Zheng et al. (2023) for the 1,466 videos in OUD Search Dataset that were not annotated by experts and train a lightweight student model on this synthetic data.
This approach minimizes API and computational costs and avoids the instability of relying on proprietary, closed-source LLMs, whose behaviors can change over time OpenAI (2024b), while achieving strong performance on our gold-standard dataset. For each myth, we train a student model, DeBERTa-v3-base He et al. (2021), with training and experimental details in §D.
Results. Table 4 shows that DeBERTa-v3-base, trained on GPT-4o-generated synthetic labels, achieves macro F1-scores between 0.60 and 0.78 across myths, with scores $\geq 0.75$ on four myths. These results demonstrate strong performance despite the models’ smaller size (186M parameters) relative to GPT-4o and its usage of synthetic data, validating both the quality of the student model and the GPT-4o-generated labels. The results highlight the models’ effectiveness on a high-stake, complex video classification task and their suitability for large-scale labeling. Full evaluation metrics are in Table 15, with additional details in §D.3.
@@ 5.3 MythTriage Implementation
With GPT-4o offering stronger performance and the distilled model enabling efficient large-scale labeling, MythTriage can combine their strengths—using the lightweight model for routine cases and deferring harder ones to the strong, but costly LLM to optimize both cost and performance. To decide which examples to defer, we use two strategies: (1) Maximum Softmax Probability (MSP), which uses the predicted class softmax probability as a simple, effective proxy for model confidence Lakshminarayanan et al. (2017), deferring examples below a chosen MSP threshold, and (2) Validation Error Tendencies (VET), which defers predictions from classes with low validation performance (e.g., class-specific F1 < 0.8). We also evaluated combining MSP and VET, deferring an example to the high-performing GPT-4o if either the MSP or VET condition is met. We discuss alternative deferral strategies, threshold selection method, and results motivating MSP and VET in §E.
@@ 7.1 Prevalence of Myths in Search Results
Kratom has the highest prevalence of myths across topics (Figure 2), with 36% of search results supporting and only 22% opposing OUD-related myths. This is concerning, given the widespread but debunked claims about Kratom’s effectiveness in treating OUD. The high prevalence of myth-supporting content may mislead users towards unsafe alternatives Mayo Clinic (2024), undermining evidence-based treatments like MAT. While Heroin and Methadone have slightly negative bias scores (-0.02 and -0.04), they still show high levels of myth-supporting content (22% and 36%). These findings can help platforms prioritize moderation on high-risk topics and inform public health officials where misinformation is most concentrated.
@@
Switching from the default “Relevance” filter increases exposure to myths. In Figure 3, the “Relevance” filter consistently returns search results that lean towards opposing myths (bias scores from -0.04 to -0.26). In contrast, alternative filters—“Upload Date,” View Count,” and “Rating”—consistently shift bias scores in a more positive direction, indicating increased prevalence of myth. This finding aligns with prior work in the domain of COVID-19 misinformation Jung et al. (2025), and suggests that users seeking recent, most-viewed, or highly-rated videos (more likes than dislikes) are more likely to encounter myths. This is concerning, as users may place more trust in popular content, and those seeking the latest information about OUD may be at greater risk of encountering myths. These findings present an opportunity for YouTube to improve moderation by enhancing safeguards in non-default search filters.